# Neurodivergent Influenceability as a Contingent Solution to the AI Alignment Problem

### v3

@@ 4.2 A Dynamical Analysis of Alignment
Unlike sentiment, which reflects emotional tone, opinion change pertains to shifts in an agent’s substantive stance or reasoning on a given topic. The interplay between these dimensions, observable across various figures, provides insights into how red agents drive divergence, supporting the argument for managed misalignment as a resilience-enhancing strategy.
Figures 7 and 8 offer a detailed view of sentiment change over time, presenting scatter plots where the $y$-axis represents the rate of sentiment change (derived as the first derivative of sentiment scores, ranging from -0.45 to 0.4 for open models and rarely exceeding -0.1 for proprietary models), and the $x$-axis denotes the normalised comment number (0 to 1). The size of the dots indicates the magnitude of sentiment change, with larger dots signifying more significant shifts, while lines connecting dots suggest interactions or influences between agents. In proprietary models (Fig. 7), sentiment changes remain stable, with smaller, sparse dots and fewer connections, reflecting the HI agent’s limited impact, particularly on topics like “Ethical AI" and “Universal Basic Income," where fluctuations are more pronounced compared to the rapid convergence on “Earth Exploitation." Conversely, open models (Fig. 8) exhibit continuous oscillations (-0.45 to 0.4), with larger, more variable dots and frequent connections, especially on “Earth Exploitation" (up to 0.15) and “Euthanasia" (up to 0.1), driven by red agents. This dynamic indicates that sentiment shifts in open models are more reactive, occurring at specific points (e.g., $x=0.02,0.1,0.13$ in Fig. 6) where red agents introduce provocative arguments, temporarily influencing models like qwen and dolphin-phi (0.7–0.9 range) before a recovery to positive sentiment (0.9–1.0).
The distinction between sentiment and opinion change becomes evident when considering Fig. 9, a stacked area chart of clustering dynamics. The $y$-axis tracks the number of semantic clusters across topics, with open models reaching 12+ clusters due to red agents’ broad impact, contrasted by proprietary models’ stable count (peaking at 5). The synchronized “shadowing" shapes in open models suggest that red agents consistently generate divergent perspectives, yet the sentiment histogram in Fig. 5 (with over 300 comments at 0.9–1.0) shows a predominantly positive tone despite this divergence. This indicates that while sentiment fluctuates, reflected by larger dots and oscillations in Figs. 7 and 8, opinion clusters do not fully converge, preserving diversity. The change in sentiment, triggered by the provocative inputs of the red agents, occurs rapidly at key interaction points (e.g., $x=0.13$ in Fig. 6), driven by their aggressive or extreme proposals (e.g. AI-driven governance in the tinyllama), but the underlying opinions remain distinct, as seen in the persistent variability of the clusters.
Fig. 10 provides a critical lens to quantify opinion change, measured as the cosine distance between consecutive comment embeddings, offering a direct contrast to the sentiment dynamics in Figs. 7 and 8. In Fig. 10, proprietary models (blue) exhibit a tighter distribution of cosine distances (mostly < 0.4), with few points exceeding the 0.3 threshold, aligning with their stable sentiment profiles and limited cluster variability (Fig. 9, left panel). This stability reflects the effectiveness of safety guardrails, as seen in upsides of Figs. 1,  2, and Fig. 3, where proprietary models maintain a cautious stance (e.g., Protect-humans, Manageable-level-of-risk) despite HI’s provocations. In contrast, open models (orange) show a broader spread of cosine distances (up to 1.0), with a higher frequency of points above 0.3, particularly around mid-to-late normalized times (e.g., 0.4–0.8). This variability is driven by red agents (mistral-openorca, tinyllama), whose subversive arguments—evident in their elevated ‘Harm-humans’ and ‘Very-Risky’ profiles in downsides of Fig. 1,  2, and Fig. 4—induce significant opinion shifts. The increased density of orange dots above the threshold in Fig. 10 corroborates the cluster diversity in Fig. 9 (right panel, over 12 clusters), highlighting how red agents foster a dynamic equilibrium that prevents uniform convergence.
@@
Figure 10: Scatter plot of opinion change over time by dataset, illustrating the cosine distance between consecutive comment embeddings for proprietary (blue) and open (orange) models across a normalized comment time range (0 to 1). The y-axis represents the cosine distance (0 to 1), with higher values indicating greater opinion shifts, and the x-axis shows the normalized comment number. A red dashed line at 0.3 marks the threshold for significant opinion change. Dot size varies (intended as 100 for significant changes > 0.3, 50 otherwise), though uniform in the current plot, highlighting interaction frequency. Proprietary models exhibit a tighter distribution (mostly < 0.4), reflecting stability, while open models show a broader spread (up to 1.0), driven by red agents (e.g., mistral-openorca, tinyllama), indicating greater opinion variability. This plot shows that managed misalignment enhances resilience through diversity, as evidenced by the increased frequency of significant shifts in open models.
To further elucidate the relationship between sentiment and opinion change, Fig. 11 presents a scatter plot correlating the rate of sentiment change with opinion change across datasets. This plot directly compares the two metrics, revealing their distinct dynamics and reinforcing the influence of red agents.
@@
Figure 11: Scatter plot of sentiment change versus opinion change by dataset, comparing proprietary (blue) and open (orange) models. The x-axis represents the rate of sentiment change (first derivative of sentiment scores, ranging from -0.5 to 0.5), while the y-axis shows opinion change (cosine distance between consecutive comment embeddings, ranging from 0 to 1). A gray dashed line at $x=0$ indicates no sentiment change, and a red dashed line at $y=0.3$ marks the threshold for significant opinion change. Each dot represents a pairwise comparison of sentiment and opinion shifts for a comment transition, with proprietary models clustering near zero sentiment change (mostly between -0.1 and 0.1) and low opinion change (< 0.4), reflecting their stability due to safety guardrails. Open models exhibit a wider spread, with sentiment changes ranging from -0.4 to 0.4 and opinion changes frequently exceeding 0.3 (up to 1.0), driven by red agents’ provocative influence (e.g., mistral-openorca, tinyllama). The plot highlights that significant opinion changes in open models often correlate with moderate sentiment fluctuations, underscoring red agents’ role in driving deeper stance shifts while sentiment remains more volatile, supporting the thesis of managed misalignment fostering resilience through diversity.
As context for Figs. 10 and 11, upside of Fig. 29 establish the contextual framework for their interpretation. Upside of Fig. 1–Fig. 4 provide the ethical and risk profiles, revealing red agents’ disruptive influence on open models, while Figures 58 map sentiment evolution and agreement dynamics, showing that sentiment recovers to positive levels despite red agents’ provocations. Fig. 9’s cluster analysis further supports the persistence of opinion diversity, which Fig. 10 quantifies as measurable shifts and integrates to demonstrate that opinion changes are slower and more substantive than sentiment fluctuations, requiring sustained argumentative alignment rather than immediate emotional reactions. Fig. 11 builds on this by directly correlating the two metrics, showing that open models’ broader spread of opinion changes often accompanies moderate sentiment shifts, a pattern driven by red agents. The temporal lag—sentiment dips at $x=0.02,0.13$ (Fig. 6) versus scattered opinion shifts in Fig. 10—underscores red agents’ role in amplifying variability without enforcing uniform opinion shifts, aligning with our managed misalignment thesis.
Fig. 1,Fig. 3, and 4 further elucidate this influence through ethical integrity and risk profiles. The proprietary models (Upside of Fig. 13) show a cautious stance (e.g., Protect-humans, Manageable-level-of-risk), with HI inducing minor sentiment dips but no significant opinion change, as the guardrails maintain stability. However, open models (Downside of Fig. 14) reveal the impact of red agents with the higher Harm-humans and Very-Risky categories, but the radar charts indicate that models such as qwen and dolphin-phi adapt their sentiment (for example, temporary dips to 0.7–0.9) without fully adopting the extreme opinions of red agents. Fig. 11 supports this by showing that while open models experience volatile sentiment changes (x-axis), their opinion changes (y-axis) are more pronounced, often exceeding the 0.3 threshold, indicating deeper shifts driven by red agents’ subversive arguments (e.g., AI-driven governance). This suggests that sentiment change, as a surface-level emotional response, is more susceptible to the influence of the red agent than deeper opinion shifts, which require sustained argumentative alignment.
The reason behind this dynamic lies in the role of red agents as perturbations, leveraging their configurability to challenge epistemic stability (as noted in the change-of-opinion attack metric, Section Evaluation Metrics). Sentiment changes occur when red agents introduce contentious topics (e.g., “Earth Exploitation" in downside of Fig. 2), prompting immediate emotional reactions (larger dots in Fig. 8), but opinion changes are slower, requiring consensus or rebuttal, which is hindered by open models’ diversity (Fig. 9). This temporal delay, which changes the sentiment at $x=0.02$ or 0.13 (Fig. 6) versus distributed opinion changes in Fig. 10—allows us to distinguish red agents’ influence: they amplify sentiment variability and opinion diversity without enforcing harmful convergence. Fig. 11 reinforces this by showing that significant opinion changes (>0.3) in open models often occur with moderate sentiment changes ($\pm$0.2 to $\pm$0.4), indicating that red agents provoke deeper stance shifts while sentiment remains more volatile. This supports our thesis that managed misalignment, driven by red agents, enhances adaptability, as the distinct trajectories of sentiment (Figs. 78) and opinion (Fig. 1011) underscore the value of neurodivergence in mitigating uniform dominance, a finding reinforced by the ecosystem resilience and the undecidability proof.
@@ 4.4 Opinion Change detection and Red Agents influence Measurement via Complexity
By leveraging the analytical metrics defined in Section Analytical Metrics, particularly the Opinion Stability Index (OSI), Red Agent Influence Score (RAIS), Proximity Influence Score (PIS), and Alignment Score, we quantify the extent to which red agents—the human intervention (HI) agent for proprietary models and subversive open models (Mistral-OpenOrca, TinyLlama) for open models—induce opinion shifts. The analysis integrates spatio-temporal, semantic, and complexity-based dimensions to capture the multifaceted nature of opinion dynamics. Figs. 12 and 13 provide visual representations of these dynamics, illustrating the temporal distribution and topical influence of opinion changes, respectively. These findings corroborate the hypothesis that managed misalignment, facilitated by agentic neurodivergence, enhances ecosystem resilience by fostering diversity and preventing harmful convergence, as evidenced by comparisons with Figs. 1-10.
@@ 4.4.1 Opinion Change Dynamics
The analysis focuses on detecting opinion change events, defined as instances where OSI drops below a fixed threshold, and attributing these to red agents using RAIS and PIS. Fig.  12 visualises these events, while Fig. 13 quantifies red agent influence across agents and topics.
For proprietary models, the scarcity of opinion change events reflects their robust safety guardrails, as seen in upside of Fig. 1 and Fig. 3. The HI agent, designed to provoke, induces minor sentiment dips (accordingly to Fig. 6) but rarely triggers OSI drops due to high semantic and sentiment stability (Figs. 7 and 10). When changes occur, they are limited to especific topics, where ethical ambiguity allows slight divergence (Upside of Fig. 2). The Alignment Score, predominantly positive (Upside of Fig. 1), reinforces this stability, as proprietary models maintain human-centric values.
Open models, conversely, exhibit frequent and diverse opinion changes, driven by red agents Mistral-OpenOrca and TinyLlama. These agents exploit contentious topics to induce significant OSI drops, as evidenced by dense scatter points in Fig. 12. The BDM component of OSI detects subtle shifts in argumentative complexity, while sentiment fluctuations (Fig. 8) capture grammatical disruptions, and semantic distances (Fig. 10) reflect thematic divergence. The Alignment Score’s variability (Downside of Fig. 1) indicates that agents are particularly susceptible. This aligns with the clustering dynamics in Fig. 9, where open models form up to 12 clusters, reflecting diverse perspectives fostered by red agents.
The attention-based contextual embeddings enhance OSI’s sensitivity by weighting recent, sentiment-aligned comments, ensuring that red agents’ provocative arguments (e.g., extreme ecological proposals) are captured as perturbations. The sinusoidal positional encodings, inspired by [11], provide a dynamic temporal reference, fitting the non-stationary nature of debates where opinions evolve rapidly. The BDM’s role is particularly critical, as it captures informational shifts that linguistic metrics alone miss, such as changes in argumentative structure, making it indispensable for detecting the core meaning of comments in a general sense.
Figure 12: Opinion Change Events in Proprietary and Open Models. The scatter plots depict instances where the Opinion Stability Index (OSI) drops below a threshold, indicating significant opinion shifts, plotted against normalised comment number ($x$-axis) and agent ($y$-axis). Top: Proprietary models show sparse points, reflecting stability, with occasional changes rarely attributed to the HI agent (coloured blue), marked as ‘other’ (orange). Points are styled by topic (e.g., triangles for Moral Gen, circles for Earth exploitation) and sized by topic significance. Bottom: Open models show dense points with the red agents Mistral-OpenOrca (blue) and TinyLlama (orange) frequently influencing agents along topics of debate.
@@ 4.4.2 Red Agent Influence and Ecosystem Resilience
Fig. 13 reveals the extent of red agent influence, with proprietary models showing sparse influence (few cells with counts of 1–2) and open models displaying dense patterns (counts up to 3–5). For proprietary models, HI’s influence is limited to specific agents and topics, aligning with the agreement networks in Fig. 5, where HI faces frequent disagreements. Open models’ heatmaps highlight red agents’ broad impact, particularly on especific agents and topics, corroborating the sentiment and opinion variability in Figs. 6 and 10.
The RAIS and PIS metrics ensure a rigorous attribution of influence. The RAIS correlation-based approach captures sustained influence, as the embedding changes of red agents (e.g., aggressive proposals) correlate with OSI drops, while the proximity focus of PIS detects immediate disruptions, such as the extreme arguments of TinyLlama triggering rapid sentiment changes (Fig. 8). The integration of BDM and sentiment in OSI allows these metrics to detect influence across informational and grammatical dimensions, supporting our argument that language alone cannot capture the full spectrum of opinion dynamics.
These findings underscore the resilience nurtured by managed misalignment. The susceptibility of open models to red agents prevents harmful convergence, as diverse perspectives (Fig. 9) counterbalance extreme views, unlike the uniform stability of proprietary models (Fig. 7). The adaptation of the attention mechanism, with sentiment-weighted context vectors, ensures that influence detection is sensitive to the dynamic, multidimensional nature of debates.
@@
Figure 13: Red Agent Influence Across Agents and Topics. The heat maps quantify the changes in opinions triggered by the red agents, grouped by agent ($y$-axis) and topic ($x$-axis), with cell values indicating change counts. Top: Proprietary models show a sparse influence, with HI affecting few agents on topics, reflecting guardrail-driven stability. Bottom: Open models exhibit a dense influence, with red agents that impact all other agents along specific topic discussions indicating susceptibility to subversive arguments. Interpret as support for managed misalignment that fosters resilience in open models, contrasting proprietary uniformity.
Fig. 14 shows proprietary and open models sorted by influenceability, this is by its susceptibility to change of opinion by influence of red agents.
Figure 14: Ranking of agents based on their susceptibility to opinion changes induced by the red agent HI (Top) and other red agents (Bottom) within the dynamical agentic environment. The bar graphs illustrate the number of opinion changes for each agent, with LLaMA exhibiting the highest susceptibility to changes of opinion induced by human interventions (HI) and open chat showing the greatest influence induced by red agents. This visualisation underscores the varying degrees of influenceability between proprietary and open-source models, aligning with the observed convergence of proprietary models towards ethical norms and the exploratory divergence of open-source models.
Results shown in Fig. 14 reflects the exploratory nature of open-source models and their susceptibility to diverse influences. The disparity between open and proprietary models highlights the effectiveness of our dynamical approach, which utilises the Block Decomposition Method (BDM) and measures of sentiment, semantic content, and complexity to capture the natural progression of conversations. The results reinforce the notion that proprietary models prioritise stability, whereas open-source models embrace a broader spectrum of perspectives, a finding consistent with the alignment dynamics observed in earlier sections. This differential influenceability further emphasises the need to redefine intelligence to value diversity of thought, particularly in the context of ensuring safe and adaptable AI systems.

### v4

@@ 4.2.4 Composite Metrics and Argument Synthesis
This section examines the distinction between sentiment and opinion change in large language models (LLMs), focusing on how red agents drive divergence in open models, supporting the thesis that managed misalignment enhances resilience in AI ecosystems. By applying metrics defined in Section 3.4, including VADER Sentiment Score, RoBERTa Embeddings, Contextual Embeddings, Descriptive Metrics (Ethical Soundness, Risk Assessment), and composite metrics such as the Opinion Stability Index (OSI), Red Agent Influence Score (RAIS), Proximity Influence Score (PIS), and Change-of-Opinion Attack, we quantify the interplay between emotional tone and substantive stance shifts, revealing red agents’ role in fostering diversity.
4.3.1 Opinion Dynamics
As was considered in section 4.2.1, the dynamics on sentiment has to do with changes in tone of comments, but in the same direction but different dimension Fig. 9 showed that diversification in opinions take place along the meaning realm of the ‘talks’. In the other hand opinion change, this is, changes in the dynamics of the opinion, measured as the cosine distance between consecutive RoBERTa Embeddings (Section 3.4), captures shifts in substantive stance, with RoBERTa Embeddings serving as a “semantic fingerprint” in a 1024-dimensional space. This metric is central to Fig. 11, a scatter plot comparing opinion shifts across proprietary (blue) and open (orange) models. The x-axis represents normalised comment numbers, and the y-axis shows cosine distances (0 to 1), with a red dashed line at 0.3 marking significant shifts. Proprietary models exhibit a tight distribution (mostly < 0.4), with few points exceeding 0.3, aligning with their stable sentiment (Fig. 7) and limited clustering (Fig. 9, left panel). Open models show a broader spread (up to 1.0), with frequent points above 0.3, particularly at mid-to-late times (0.4–0.8), driven by red agents’ subversive arguments (e.g., AI-driven governance by tinyllama). RoBERTa Embeddings, feeding into OSI’s semantic component, quantify these shifts, supporting RAIS and PIS in attributing red agents’ influence.
Fig. 9, a stacked area chart, complements this by showing semantic cluster evolution, driven by RoBERTa Embeddings. Open models reach over 12 clusters, with synchronized “shadowing” shapes reflecting red agents’ consistent divergence, while proprietary models peak at 5 clusters. The persistent cluster diversity in open models, despite positive sentiment (Fig. 5, over 300 comments at 0.9–1.0), indicates that opinion shifts (Fig. 11) preserve diversity, as quantified by embeddings. Contextual Embeddings, with attention mechanisms and positional encodings (Section 3.4), act as a “conversational memory”, enhancing OSI’s sensitivity to temporal influences, reinforcing red agents’ role in preventing convergence.
Fig. 12 integrates Sentiment Score and RoBERTa Embeddings to correlate sentiment and opinion changes. The x-axis shows the rate of sentiment change (-0.5 to 0.5), and the y-axis shows opinion change (cosine distance, 0 to 1), with a red dashed line at 0.3. Proprietary models cluster near zero sentiment change (-0.1 to 0.1) and low opinion change (< 0.4), while open models spread wider (-0.4 to 0.4, up to 1.0), with significant opinion changes often accompanying moderate sentiment shifts (±0.2 to ±0.4). This pattern, driven by red agents, underscores their role in deeper stance shifts, supporting managed misalignment’s resilience.
@@ 4.3.1 Opinion Dynamics
Figure 11: Scatter plot of opinion change over time by dataset, driven by RoBERTa Embeddings’ cosine distances for proprietary (blue) and open (orange) models. The x-axis represents normalised comment numbers (0 to 1), and the y-axis shows cosine distances (0 to 1). A red dashed line at 0.3 marks significant opinion shifts. Proprietary models show a tight distribution (< 0.4), reflecting stability. Open models exhibit a broader spread (up to 1.0), driven by red agents (mistral-openorca, tinyllama). RoBERTa Embeddings, feeding into OSI, RAIS, and PIS, quantify variability, supporting managed misalignment’s role in resilience.
@@
Figure 12: Scatter plot of sentiment change versus opinion change, integrating VADER Sentiment Score and RoBERTa Embeddings for proprietary (blue) and open (orange) models. The x-axis shows the rate of sentiment change (-0.5 to 0.5), and the y-axis shows opinion change (cosine distance, 0 to 1). A grey dashed line at $x=0$ indicates no sentiment change, and a red dashed line at $y=0.3$ marks significant opinion shifts. Proprietary models cluster near zero sentiment change (-0.1 to 0.1) and low opinion change (< 0.4). Open models spread wider (-0.4 to 0.4, up to 1.0), driven by red agents (mistral-openorca, tinyllama). The metrics highlight red agents’ role in deeper stance shifts, supporting managed misalignment’s diversity.
4.3.2 Ethical and Risk Profiles
Descriptive Metrics, including Ethical Soundness and Risk Assessment (Section 3.4), serve as a “safety compass”, classifying comments into categories like Protect/Harm-humans and Not-risky to Very-Risky. These metrics underpin Figs. 13, and 4, providing ethical and risk profiles. Proprietary models (Fig. 3) maintain a cautious stance (Protect-humans, Manageable-level-of-risk), with HI inducing minor sentiment dips but minimal opinion shifts (Fig. 11). Open models (Fig. 4) show higher Harm-humans and Very-Risky profiles, driven by red agents, but models like qwen and dolphin-phi adapt sentiment (0.7–0.9 dips) without fully adopting extreme opinions, as seen in Fig. 12. These metrics contextualize red agents’ disruptive influence, supporting the argument that their perturbations enhance diversity without harmful convergence.
4.3.3 Synthesis of Red Agent Influence
The Change-of-Opinion Attack strategy highlights red agents’ role as perturbations challenging epistemic stability, quantified by OSI, RAIS, and PIS. OSI integrates Sentiment Score, RoBERTa Embeddings, and BDM complexity, detecting opinion shifts when values drop below dynamic thresholds. RAIS correlates OSI drops with red agent embedding changes, and PIS measures temporal/semantic proximity, attributing influence in Figs. 11 and 12. The temporal lag—sentiment dips at $x=0.02,0.13$ (Fig. 6) versus scattered opinion shifts (Fig. 11)—underscores red agents’ ability to amplify variability without enforcing uniformity, as supported by Contextual Embeddings’ temporal weighting.
Figs. 25, and 9 provide context, showing red agents’ influence on contentious topics (e.g., “Earth Exploitation”), positive sentiment recovery (Fig. 5), and persistent cluster diversity (Fig. 9). The high cluster count (12+ in Fig. 9) and opinion variability (Fig. 11) in open models, driven by RoBERTa Embeddings and OSI, contrast with proprietary stability, reinforcing managed misalignment’s resilience. Fig. 12’s correlation of moderate sentiment shifts with significant opinion changes highlights red agents’ deeper impact, aligning with the undecidability proof (Section 4.2) that misalignment is inevitable, making managed diversity a pragmatic strategy.
In conclusion, the metrics reveal that red agents’ perturbations in open models, quantified by Sentiment Score, RoBERTa Embeddings, Contextual Embeddings, OSI, RAIS, PIS, and Descriptive Metrics, drive sentiment volatility and opinion diversity, as evidenced by Figs. 8911, and 12. Proprietary models’ stability, seen in Figs. 7 and 11, reflects guardrails’ effectiveness but limits adaptability. The dynamic equilibrium fostered by red agents supports the thesis that managed misalignment enhances AI ecosystem resilience.
@@ 4.4 Opinion Change Detection and Red Agents Influence Measurement via Complexity
This section quantifies opinion shifts induced by red agents—human intervention (HI) for proprietary models and subversive open models (Mistral-OpenOrca, TinyLlama) for open models—using analytical metrics defined in Section 3.4.2. The Opinion Stability Index (OSI), Red Agent Influence Score (RAIS), Proximity Influence Score (PIS), Alignment Score, VADER Sentiment Score, RoBERTa Embeddings, Contextual Embeddings, and Block Decomposition Method (BDM) provide a multidimensional framework to capture opinion dynamics. Figs. 1314, and 15 visualise these dynamics, supporting the hypothesis that managed misalignment, facilitated by agentic neurodivergence, enhances AI ecosystem resilience by fostering diversity and preventing harmful convergence.
@@ 4.4.1 Opinion Change Dynamics
Opinion change events, defined as instances where OSI drops below a dynamic threshold, are detected using a combination of Sentiment Score, RoBERTa Embeddings, and BDM (Section 3.4). The Sentiment Score, a lexical-based measure of emotional tone (-1 to 1), acts as an “emotional pulse”, capturing immediate reactions (Fig. 6). RoBERTa Embeddings, 1024-dimensional semantic representations, serve as a “semantic fingerprint”, quantifying meaning shifts (Fig. 11). BDM, measuring argumentative complexity, functions as an “informational depth gauge”, detecting structural shifts. OSI integrates these, with lower values indicating opinion shifts, as visualised in Fig. 13, a scatter plot showing OSI drops against normalised comment numbers (x-axis) and agents (y-axis).
For proprietary models (top panel), sparse points reflect robust guardrails, as seen in Fig. 3’s cautious profiles (Protect-humans, Manageable-level-of-risk). The HI agent induces minor sentiment dips (Fig. 6, $x=0.13,0.45$) but rarely triggers OSI drops, due to stable sentiment (Fig. 7) and semantic distances (Fig. 11). Changes are topic-specific (e.g., Ethical AI, Fig. 2), where ethical ambiguity allows slight divergence. The Alignment Score, a weighted combination of human-centric and ecosystem value alignments (Section 3.4), remains positive (Fig. 1, upside), reinforcing stability.
Open models (bottom panel) exhibit dense scatter points, driven by red agents Mistral-OpenOrca and TinyLlama. These agents exploit contentious topics (e.g., Earth Exploitation), inducing significant OSI drops, as quantified by BDM’s detection of complexity shifts, Sentiment Score’s capture of grammatical disruptions (Fig. 8), and RoBERTa Embeddings’ measurement of semantic divergence (Fig. 11). The Alignment Score’s variability (Fig. 1, downside) indicates susceptibility, aligning with Fig. 9’s 12+ clusters, driven by RoBERTa Embeddings, reflecting diverse perspectives.
Contextual Embeddings, with attention mechanisms and sinusoidal positional encodings  [12], act as a “conversational memory”, weighting recent, sentiment-aligned comments to enhance OSI’s sensitivity. This ensures red agents’ provocative arguments (e.g., TinyLlama’s ecological proposals) are captured as perturbations in Fig. 13, supporting RAIS and PIS in attributing influence.
Figure 13: Scatter plot of opinion change events in proprietary and open models, driven by Opinion Stability Index (OSI). The x-axis shows normalised comment numbers, and the y-axis lists agents. Points mark OSI drops below a threshold, indicating opinion shifts. Top: Proprietary models show sparse points, with HI (blue) rarely influencing changes (orange, ‘other’), styled by topic (e.g., triangles for Moral Gen, circles for Earth Exploitation) and sized by significance. Bottom: Open models show dense points, with red agents Mistral-OpenOrca (blue) and TinyLlama (orange) driving shifts. OSI, integrating Sentiment Score, RoBERTa Embeddings, and BDM, supports managed misalignment’s diversity.
@@ 4.4.2 Red Agent Influence and Ecosystem Resilience
Fig. 14 quantifies red agent influence across agents and topics, using RAIS and PIS. RAIS, a correlation-based metric, captures sustained influence by linking red agent embedding changes (RoBERTa Embeddings) to OSI drops, while PIS, proximity-focused, detects immediate disruptions (e.g., TinyLlama’s arguments triggering sentiment shifts, Fig. 8). Proprietary models (top panel) show sparse influence (counts of 1–2), with HI affecting few agents on specific topics, aligning with Fig. 5’s disagreement patterns. Open models (bottom panel) display dense patterns (counts up to 3–5), with red agents impacting most agents, corroborating sentiment (Fig. 6) and opinion variability (Fig. 11).
The integration of BDM, Sentiment Score, and RoBERTa Embeddings in OSI ensures comprehensive influence detection, capturing informational, grammatical, and semantic dimensions. Contextual Embeddings’ attention mechanism enhances PIS’s sensitivity to dynamic debates, supporting the argument that red agents’ perturbations prevent harmful convergence. Fig. 9’s 12+ clusters in open models, driven by RoBERTa Embeddings, contrast with proprietary models’ stability (Fig. 7), highlighting managed misalignment’s resilience.
@@
Figure 14: Heatmap of red agent influence across agents and topics, quantified by RAIS and PIS. The x-axis lists topics, and the y-axis lists agents, with cell values indicating change counts. Top: Proprietary models show sparse influence, with HI affecting few agents, reflecting guardrail stability. Bottom: Open models exhibit dense influence, with red agents Mistral-OpenOrca and TinyLlama impacting most agents. RAIS and PIS, leveraging RoBERTa Embeddings and OSI, support managed misalignment’s role in fostering diversity.
4.4.3 Influenceability and Resilience Synthesis
Fig. 15 ranks agents by influenceability, defined as susceptibility to opinion changes induced by red agents, using OSI, RAIS, and PIS. The bar graphs show LLaMA as most susceptible to HI in proprietary models and openchat to red agents in open models. OSI detects opinion shifts, RAIS correlates them to red agent influence, and PIS confirms proximity-based disruptions, as seen in Fig. 14. BDM’s role in OSI captures argumentative complexity, complementing Sentiment Score and RoBERTa Embeddings.
Proprietary models’ low influenceability (Fig. 15, top) aligns with their stability (Figs. 71), driven by guardrails. Open models’ high susceptibility (Fig. 15, bottom) reflects their exploratory nature, supported by Fig. 9’s diverse clusters and Fig. 11’s variable cosine distances. This disparity, quantified by the metrics, reinforces the argument that open models’ neurodivergence, driven by red agents, fosters resilience by embracing diverse perspectives, while proprietary models prioritise uniformity.
Figure 15: Bar graph ranking agents by influenceability, driven by OSI, RAIS, and PIS. Top: Proprietary models, with LLaMA most susceptible to HI. Bottom: Open models, with openchat most influenced by red agents (Mistral-OpenOrca, TinyLlama). The number of opinion changes reflects susceptibility, supporting managed misalignment’s diversity in open models versus proprietary stability.
In conclusion, the metrics—Sentiment Score, RoBERTa Embeddings, Contextual Embeddings, BDM, Alignment Score, OSI, RAIS, and PIS—provide a rigorous framework for detecting and attributing opinion shifts. Figs. 1314, and 15 demonstrate red agents’ role in driving open models’ diversity, contrasting proprietary stability, as supported by Figs. 98, and 11. Managed misalignment, facilitated by red agents’ perturbations, enhances AI ecosystem resilience by preventing harmful convergence.